OpenStudy (anonymous):

PLz help!! Andrew has two part time jobs. One pays $6 per hour and the other pays $10 per hour. He wants to make $366 per week. Write an equation in standard form that describes this situation. If he is only allowed to work 15 hours per week at the $10 per hour job, how many hours does he need to work per week in his $6 per hour job in order to achieve his goal

OpenStudy (anonymous):

I don't understand how to write this out?

OpenStudy (the_fizicx99):

"standard form" ?

OpenStudy (anonymous):

?

OpenStudy (the_fizicx99):

I'm guessing Ax + By = C?

random231 (random231):

ohh wat i wrote was right! 6*x+10*y=366 x and y are hours per week on each rate!!

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

1) 10x + 6y = 366 2) 10* 15 + 6y = 366 -> 6y = 216 -> y = ...

OpenStudy (the_fizicx99):

Then it should be: 10 * 15 = 150, 366 - 150 = 216/6 = 36 10x + 6y = 366 x = 15, y = 36 ?

OpenStudy (anonymous):

thanks guys that makes more sense then it did :)

OpenStudy (the_fizicx99):

10x + 6y = 366 x + y = 51